A comparative balance sheet for Flint Corporation is presented as follows. Assets Cash Accounts receivable Inventory Land EquPrepare a statement of cash flows for 2020 for Flint Corporation. (Show amounts that decrease cash flow with either a-sign e.Issued common stock to retire $ of bonds outstandingDetermine Flint Corporations current cash debt coverage, cash debt coverage, and free cash flow. (Round current cash debt co
Answer #1
Cash Flow Statement
Indirect Method
Cash Flow from Operating Activities
Net Income $           127,920
Adjustments
Depreciation $              26,810 =70460-43650
Changes in Current Assets/ Current Liabilities
Increase in Accounts Receivable $            (15,810) =67650-83460
Decrease in inventory $                9,190 =190650-181460
Decrease in Accounts Payable $            (13,190) =35460-48650
Total Adjustments $                7,000
Cash from Operating Activities $         134,920
Cash flow from Investing Activities
Sale of Land $              39,190 =111650-72460
Purchase of Equipment $           (59,810) =201650-261460
Net cash used In investing activities $         (20,620)
Cash flow from Financing Activities
Dividend Paid $           (63,490)
Net cash used in financing activities $         (63,490)
Increase in Cash $            50,810
Opening Balance of Cash $            22,000
Closing Balance of Cash $            72,810
Non Cash Financing Activities
Issue common stock to retire $50,000 of bonds outstanding

Current Cash Debt coverage Ratio = Net Cash provided by operating activities / Average Current Liabilities
= $134920 / $42055 = 3.21
Average Current liabilities = ($35460+48650)/2 = $42055

Cash Debt coverage Ratio = Net Cash provided by operating activities / Average Total Liabilities
= $134920 / $217055 = 0.62
Average Current liabilities = ($35460+48650+150000+200000)/2 = $217055

Free Cash flow = Net Cash provided by operating activities - Net cash used in investing activities
= $134920-20620 = $114300

Flint has high liquidity. Its financial flexibility is good